Please help me with 6a calculations
Experimental Procedure, Part A.6 versus Part B.2.
Answer #1

Q5.

Vadded = Vf-Vi = 13.9-1.7 = 12.2 mL

Q6.

moles = MV

moles = (0.048)(12.2*10^-3) = 0.0005856 moles of HCl

Q7.

assume H+ = OH-

so

0.0005856 mol of H+ = 0.0005856 molf of OH-

For Q8:

[OH-]eq = [OH-]added - [OH-]reacted

[OH-]reacted = 0.0005856 mol / 25 mL = 0.023424 M

[OH-]added --> from your lab practice data

Q9.

[Ca+2] = from your lab data

Q10

Molar solubility

[Ca+2] = S

[Ca+2] = (from Q9)

S = [Ca+2]

Q11

Ksp = [Ca+2][OH-]^2

so

[Ca+2] = S

[OH-] = 2S

Ksp = S*(2S)^2) = 4*S^3

Ksp = 4*(S^3) S from Q9